Database last updated on Tue, 27 Jan 2026 17:40:45 +0000 by localhost
Critical warnings from b-files' comments for days between 2025112 and 2025112 (0 sec, 2 rows)
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
| rl: | 2025 111 05 33 59 Motor #2: appears stuck with a reference sensor activated. | 3 | 1 | B11225.224 | TORON | #224 | 05:25:52 | | 1 | -1 |
| rl: | 2025 112 18 35 38 New value set for watermark SUPPLY.VOLTAGE.LOW | 6 | 3 | B11225.224 | TORON | #224 | 05:26:01 | | 1 | -1 |
Instrument constant changes for the week before 2025112 (2 sec, 0 rows)
SL burned out? (AP results when SL is ON) (0 sec, 2 rows)
Non-critical warnings from b-files' comments for days between 2025112 and 2025112 (0 sec, 3 rows)
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
| rl: | 2025 111 05 31 16 Reset Warm -> Tepid from 2025 111 05 31 11 until 2025 111 05 31 11 . Counts: 1,0 | 1 | 3 | B11225.224 | TORON | #224 | 05:25:49 | | 1 | -1 |
| re: | Azimuth discrepancy = -177 | 1 | 99 | B11225.224 | TORON | #224 | 05:34:05 | -177 | -30 | 30 |
| ds: | Suppressed oscillation of filter wheel. | 3 | 5 | B11225.224 | TORON | #224 | 13:52:19 | | 1 | -1 |
Measurements and tests summary.
| _jday | _loc | _serial | _skc | ds | dt | ed | fz | hg | rs | sl | uv | zs |
| 112 | TORONTO | #224 | ecsc,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 93.033) | 76 | 2 | 2 | 6 | 16 | 2 | 7 | 31 | 38 |
(0 sec)
Number of days since last dome/window cleaning (0 sec, 2 rows)
| serial | days |
| 224 | more than a year |
Missing files (0 sec, 1 rows)
| Serial | last data | days ago |
| 224 | 2025113 | 0-1 |
Run/Stop results for last week. (0 sec, 1 rows)
| Serial | Total | Failed 0 | Failed dark | Failed 1 | Failed 2 | Failed 3 | Failed 4 | Failed 5 | Failed double |
| 224 | 16 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
Standard lamp intensity analysis (0 sec, 1 rows)
| Serial | min | max | filter | fixable | 30-day history |
| #224 | 2663230 | 2665860 | 0 | yes |